OpenStudy (hazelbrown):

Express (2-root 3)^2

OpenStudy (ayonnaleflore):

@hazelbrown im not understanding the question and are there any answer choices

OpenStudy (ayonnaleflore):

is it \[\sqrt[2]{3}^{2}\]

zepdrix (zepdrix):

Hey Hazel :) Wecome to OpenStudy! Do you need help expanding the binomial?

zepdrix (zepdrix):

\[\large\rm (2-\sqrt3)^2\]

OpenStudy (hazelbrown):



OpenStudy (ayonnaleflore):

well first you would start by doing everything that is in parenthesis

OpenStudy (hazelbrown):

what does parenthesis mean

OpenStudy (ayonnaleflore):

\[(2-\sqrt{3})\]

OpenStudy (ayonnaleflore):

would would need to solve this part first..... let me know what you get

zepdrix (zepdrix):

You can either distribute normally, or simply use your rule for expanding a square binomial,\[\large\rm (a-b)^2=a^2-2ab+b^2\]

OpenStudy (ayonnaleflore):

2 would be a and 3 would be b correct